A magic square is an arrangement of n2 numbers into n rows and n columns using distinct numbers from 1 up to n2 such that the sum in any row, any column or any of the two diagonals is fixed. Consider the 4 by 4 magic square below:


a1 a5 2 13a2 10 11 a7a3 6 a4 124 15 a6 1
What is the value of a1 ?

1
Expert's answer
2021-12-01T13:56:07-0500

Calculate the magic constant


"sum=n[(n^2+1)\/2]"

"sum=4[(4^2+1)\/2]=34"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 a_1 & a_5 & 2 & 13 \\ \hline
 a_2 & 10 & 11 & a_7 \\ \hline
 a_3 & 6 & a_4 & 12 \\ \hline
 4 & 15 & a_6 & 1 \\ 
\end{array}"

"a_5+10+6+15=34=>a_5=3"

"a_1+3+2+13=34=>a_1=16"

"16+10+a_4+1=34=>a_4=7"

"a_3+6+7+12=34=>a_3=9"

"16+a_2+9+4=34=>a_2=5"

"5+10+11+a_7=34=>a_7=8"

"4+15+a_6+1=34=>a_6=14"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 16 & 3 & 2 & 13 \\ \hline
 5 & 10 & 11 & 8 \\ \hline
 9 & 6 & 7 & 12 \\ \hline
 4 & 15 & 14 & 1 \\ 
\end{array}"

"a_1=16"